OpenStudy (anonymous):

Using complete sentences, explain how you would factor completely x^12 – 8.

OpenStudy (ash2326):

We have \[x^{12}-8\] we know \[x^{12}=(x^4)^3\] and\[8=2^3\] so we have now the expression as \[({x^4})^3-2^3\] we know \[(a-b)^3=(a-b)(a^2+ab+b^2)\] using this we get \[({x^4})^3-2^3=(x^4-2)(x^8+2x^4+4)\]

OpenStudy (ash2326):

or \[x^{12}-8=(x^4-2)(x^8+2x^4+4)\]